6.35 is not equivalent to:

a. 6 7/20

b. 635%

c. 635/100

d. 6 1/35


1
Expert's answer
2021-05-11T13:32:49-0400

To determine the value among the options which is not equivalent to 6.35 we test the options as follows:

(a)."6\frac{7}{20}"

"6\frac{7}{20}=\frac{127}{20}=6.35"


(b). "635\%"

"\frac{635}{100}=6.35"


(c). "\frac{635}{100}"

"\frac{635}{100}=6.35"


(d). "6\frac{1}{35}\:"

"\:\:\:\:\:6\frac{1}{35}=\frac{211}{35}=6.029"

Therefore the correct answer is (d). "6\frac{1}{35}\:"